# A robust mixed-effects quantile regression model using generalized Laplace mixtures to handle outliers and skewness

### v1

@@ Abstract
1   Introduction
Mixed-effects quantile regression models have become increasingly important in statistical analyses because they account for the correlation structure inherent in repeated or clustered measurements, thus facilitating more accurate inference of the conditional distribution across various quantiles [1, 2]. Under a standard mixed-effects framework, this approach is well-suited to heteroscedastic settings, where variability can differ substantially across individuals or groups. In longitudinal clinical trials, for example, certain treatments may have a stronger effect on patients in the lower or upper tails of the response distribution than those near the mean. By capturing these diverse responses, quantile-based methods reveal treatment effects that mean-focused analyses might overlook. The asymmetric Laplace (AL) distribution has long been the favored choice for parametric mixed-effects quantile regression because its fixed skewness parameter directly targets the quantile of interest [2].
@@ 1   Introduction
To overcome the AL distribution’s forced skewness constraint, Yan and Kottas [3] proposed the generalized AL (GAL) distribution, which gives greater flexibility for asymmetric and heavy-tailed data. Yu and Yu [4] further extended this approach to mixed-effects models in a quantile regression framework. By adding an estimable shape parameter, the GAL distribution accommodates a wider range of skewness and kurtosis, making it more adaptable than the AL in many settings.
@@
However, while the GAL distribution is more flexible, it can still be vulnerable to extreme outliers, especially in heavy-tailed data. By construction, its location parameter is the $p_{0}$th quantile. When $p_{0}$ is close to 0.5, both tails remain relatively heavy. But as $p_{0}$ moves away from 0.5 (i.e., $p_{0}\to 0^{+}$ or $p_{0}\to 1^{-}$), the skewness parameter “lightens" one tail while weighting the other more heavily, leaving the lighter-tail side more susceptible to outliers. To address this, we propose a more robust approach: the contaminated GAL (cGAL) distribution. Constructed as a mixture of two GAL components, the cGAL inflates the variance of the second component to account for extreme observations, effectively mitigating outlier sensitivity. Crucially, it retains the property that the location parameter aligns with the targeted quantile, preserving interpretability while reducing sensitivity to skewness and outliers.
@@
Other relevant work includes Barata [5], which examined dynamic quantile regression using the GAL distribution in Bayesian settings, though without extending it to mixed models or multivariate data with cluster-level random effects. Florez et al. [6] applied quantile regression to multivariate and longitudinal data with the multivariate generalized hyperbolic distribution, which effectively models complex dependence but omits random effects and relies on the Bessel function, complicating computation. Wichitaksorn et al. [7] explored flexible Bayesian quantile regression via scale mixtures of truncated normals. However, akin to AL, their methods lack asymmetry at middle quantiles and produce lighter tails at extremes, increasing outlier sensitivity. Moreover, these approaches do not extend to random effects. Bernardi et al. [8] employed the skew exponential power distribution for quantile regression, yielding more flexible tails than the AL but applying it only to fixed-effects models, thus excluding random and mixed model contexts.
@@
In this paper, we adopt a Bayesian approach to estimate parameters of the mixed-effects cGAL model with JAGS [9]. This approach enables full posterior inference and does not rely on asymptotic approximations. Full posterior inference can be especially beneficial for smaller or more complex models. Because JAGS supports flexible, user-defined model specifications, our conceptual approach can be translated directly into code, which conveniently facilitates the implementation of the proposed mixed-effects quantile regression model.
@@
2   ACTG 315 HIV clinical trial
The ACTG 315 dataset, available in the ushr package in R [10], includes longitudinal measurements of HIV viral load (log10 RNA copies/mL) over time. It features data on 46 patients, with the longest measurement recorded on Day 196 after baseline (Day 0). Models of HIV viral load dynamics typically show a biphasic decline: an initial rapid reduction in viral load due to effective antiretroviral therapy, followed by a slower stabilization phase as the virus is further suppressed [11]. Traditionally, many analyses (e.g., generalized linear models) have concentrated on the mean response, giving insights into average patient trajectories. Yet such approaches often overlook the considerable variability among patients, as shown in Figure 1. Some patients experience sharper declines, while others have irregular fluctuations or slower viral suppression, indicating outliers or subgroups with distinct response patterns [12]. A more flexible framework, such as quantile regression, is therefore needed to capture the full distribution of viral load trajectories and better understand HIV progression across diverse patient subgroups.
@@ 2   ACTG 315 HIV clinical trial
3   Laplace distributions for quantile regression
@@
3.1 Asymmetric Laplace distribution
@@
The AL distribution can also be viewed as a normal-exponential mixture:
@@
3.2 Generalized asymmetric Laplace distribution
Next, consider the GAL distribution suggested by Yan and Kottas [3], which generalizes the AL. As with the AL distribution, we fix $\mu$ to be the $p_{0}$th quantile so that
@@ 3.2 Generalized asymmetric Laplace distribution
The GAL distribution can also be expressed as a triple mixture of normal, exponential, and truncated normal variables:
@@
3.3 Contaminated generalized asymmetric Laplace distribution
To increase the robustness of a quantile model to outlying values, we introduce the cGAL distribution, formed as a mixture of two GAL distributions. Both components share the same location $\mu$ but differ in scale parameters: $\sigma$ for the main component and $\tau_{0}\sigma$ for the “contaminated" component, where $\tau_{0}>1$. Let $\alpha\in\left(0,1\right)$ be the mixture weight. The cGAL PDF is:
@@ 3.3 Contaminated generalized asymmetric Laplace distribution
By inflating the scale in the second component, the cGAL distribution can better accommodate outliers. This two-component, scale-inflated mixture approach closely resembles the well-known contaminated normal distribution [13, 14].
@@
Figure 3 presents the sample kurtosis and skewness for the GAL and cGAL ($\tau_{0}=10$, $\alpha=0.5$) distributions as functions of the parameter $\gamma$ to illustrate the impact of contamination in the cGAL distribution on both the tails and asymmetry of the distribution in comparison to the GAL model for two quantiles ($p_{0}\in\left\{0.1,0.5\right\}$). The left and right L-statistic kurtosis [15] are displayed alongside both distributions’ overall kurtosis and skewness. Details on the calculation of the L-statistic kurtosis can be found in Appendix A. For instance, the GAL distribution exhibits lighter-than-normal left tails at the lower decile ($p_{0}=0.1$). However, with the introduction of contamination in the cGAL distribution, the left tail becomes heavier, making it more robust to outliers, particularly for modeling at the lower decile ($p_{0}=0.1$).
@@
Figure 3: Kurtosis and skewness for GAL and cGAL ($\tau_{0}=10$, $\alpha=0.5$) over $\gamma$. Left and right L-statistic kurtosis [15] are shown in blue and red, respectively. A dotted black line marks the normal-reference kurtosis of $0.4142$. The overall kurtosis (green) and skewness (black) for GAL ($\tau_{0}=1$, solid) vs. cGAL ($\tau_{0}=10$, dashed) highlight cGAL’s heavier tails and increased robustness. A dotted line at 00 on the skewness plot indicates symmetry.
4   Bayesian mixed-effects quantile regression model
@@
where $\mu_{ij}\left(p_{0}\right)$ is the conditional $p_{0}^{\text{th}}$ quantile of $y_{ij}$ given the covariates and $\epsilon_{ij}\left(p_{0}\right)$ is the error term such that the $p_{0}^{\text{th}}$ quantile of $\epsilon_{ij}\left(p_{0}\right)$ is zero. Thus, the distribution of $\epsilon_{ij}\left(p_{0}\right)$ is centered at zero in the “$\tau$-quantile" sense.
@@
The mixed-effects quantile model is formulated under the three different distributional frameworks in Section 3: the AL, GAL, and cGAL distributions. In the AL model, $y_{ij}$ follows $f_{\text{AL}}\left(y_{ij}\left|\mu_{ij}\left(p_{0}\right),\sigma,p_{0}\right.\right)$. The GAL model introduces a skewness parameter, giving $f_{\text{GAL}}\left(y_{ij}\left|\mu_{ij}\left(p_{0}\right),\sigma,\gamma,p_{0}\right.\right)$. Lastly, the cGAL model incorporates an additional contamination parameter to handle outliers, leading to $f_{\text{cGAL}}\left(y_{ij}\left|\mu_{ij}\left(p_{0}\right),\sigma,\gamma,\alpha,p_{0}\right.\right)$.
In the cGAL setting, “outliers" or extreme data points are partially explained by the contaminated component.
For the fixed effects $\bm{\beta}$, we assume a multivariate normal prior with mean $\bm{0}$ and covariance matrix $1000\cdot\bm{I}_{p}$:
${\bm{\beta}}\sim\text{MVN}\left({\bm{0}},1000\cdot{\bm{I}}_{p}\right),$ (18)
@@
$\sigma\sim t_{+}\left(0,10,3\right),$ (19)
@@
indicating a $t$-distribution with 3 degrees of freedom, centered at 0, scale 10, and truncated to be positive.
For the parameter $\gamma$, we assume a uniform prior over the interval $\left(L_{\gamma},U_{\gamma}\right)$:
$\gamma\sim\text{Uniform}\left(L_{\gamma},U_{\gamma}\right).$ (20)
@@
We specify the matrix-generalized half-$t$ (MGH-$t$) prior distribution of Huang and Wand [16] for the inverse covariance matrix $\bm{\Sigma}^{-1}$ as a hierarchical mixture of a Wishart distribution and gamma distributions for the diagonal elements of a scale matrix, $\bm{\Psi}$. Specifically, the diagonal elements of $\bm{\Psi}$, denoted as $\psi_{ii}$, follow independent gamma distributions with shape and scale parameters set to $0.5$ and $1/50^{2}$, respectively: $\psi_{ii}\sim\text{Gamma}\left(0.5,1/50^{2}\right)$. The off-diagonal elements $\psi_{ij}$ for $i\neq j$ are set to zero. The inverse covariance matrix, $\bm{\Sigma}^{-1}$, follows a Wishart distribution with inverse scale matrix $4\bm{\Psi}$ and degrees of freedom equal to $d+1$. This hierarchical formulation corresponds to specifying a half-$t$ prior for the standard deviation terms of $\bm{\Sigma}$ and a uniform prior, $\text{Uniform}\left(-1,1\right)$, for the correlation parameters.
We assign the contamination weight $\alpha$ a weakly informative $\text{Beta}\left(1,9\right)$ prior, where the shape parameters 1 and 9 yield a mean of 0.1 and a heavily right-skewed distribution. This choice reflects our expectation that only a small fraction of observations arise from the “contaminated” component while still allowing the data to support larger values of $\alpha$ if necessary. If $\alpha$ grows too large (close to 1), the second (inflated-scale) component ceases to be a “contamination” and instead dominates the mixture. We also fix $\tau_{0}=10$ to provide a sufficiently large “scale” inflation, capturing heavy-tailed behavior without overwhelming the main component. Under these settings, the model remains close to a single-component GAL distribution but retains the flexibility to accommodate outliers through the contamination mixture.
5   Application to HIV dataset
5.1 Nonlinear mixed-effects model
@@
The NLME model we explore is a specific case of the general mixed-effects quantile regression framework introduced in Section 4, tailored for modeling the HIV viral load in the ACTG 315 dataset. In this case, the model is nonlinear, specified as a biphasic (two-phase) exponential decay process that captures both short-term and long-term dynamics of HIV viral load [11]. We model the observed viral load for patient $i$ at time $t_{ij}$ as a combination of two exponential decay processes to explain the biphasic viral decline. The model accounts for both short-term and long-term decay components and includes random effects to capture patient-specific variability.
We model the observed viral load $y_{ij}$ as follows:
@@
The random effects vector $\bm{b}_{i}=\left(b_{i1},b_{i2},b_{i3},b_{i4}\right)^{\prime}$ is assumed to follow a multivariate normal distribution as in Equation (17).
@@
The choice of $\tau_{0}=10$ for the cGAL model corresponds to 10-fold contamination in the scale parameter $\sigma$ to account for potential outliers in the viral load data. While we fix $\tau_{0}$ here for simplicity, one may extend the model to treat $\tau_{0}$ as another parameter to be estimated. However, this would increase the posterior’s complexity and potentially require careful identifiability considerations.
@@
The above NLME model, based on the GAL distribution, was also applied to the ACTG 315 dataset by Yu and Yu [4]. However, they implemented a custom sampler instead of JAGS and adopted different prior specifications. For instance, they employed a Wishart prior for $\bm{\Sigma}$, whereas we use the matrix-generalized half-$t$ distribution.
5.2 Model implementation, adequacy, and comparison
@@
All models were implemented and ran in JAGS [9] through the runjags package in R [17]. The Bayesian models were executed using Markov chain Monte Carlo (MCMC) sampling, with multiple chains initiated to ensure a thorough exploration of the posterior distribution. Convergence diagnostics were checked for each model using the potential scale reduction factor ($\hat{R}$), with values below 1.05 indicating acceptable convergence across chains.
For the AL and GAL models, we utilized their mixture representations in Equations (5) and (11) to streamline the MCMC sampling process. In contrast, for the cGAL model, we used the “ones" trick to define its custom likelihood in JAGS. This method uses binary indicator variables that follow a Bernoulli distribution. The probabilities are proportional to the likelihood, scaled by a constant to ensure they remain valid. Compared to the AL and GAL models’ mixture representation, this approach increases computational demands.
@@
The Kullback-Leibler (K-L) divergence is employed to assess the adequacy of the three models to identify influential observations and outliers [18, 19], as outlined in Appendix B.
We conducted a simulation-based residual analysis using the R package DHARMa [20, 21] to evaluate the adequacy of the three models. This method compares the observed data to the posterior predictive distribution, conditioned on the random effects for each observation, thereby generating scaled residuals between 0 and 1. We used several tests provided by DHARMa to assess these scaled residuals statistically. The Kolmogorov-Smirnov test evaluated the uniformity of the residual distribution; a parametric dispersion test checked for overdispersion or underdispersion, and a binomial test flagged observations significantly deviating from model expectations. We report p-values from these tests to determine each model’s adequacy of fit.
The loo package [22] in R was used to perform leave-one-out (LOO) cross-validation to compare the models’ predictive performance. To ensure numerical stability in the computation of the GAL and cGAL density functions, especially when dealing with exponentials of large magnitudes, we adopted a numerically stable implementation of the log-density function. The original GAL density function involves exponential and CDF terms that can cause overflow or underflow issues. To mitigate this, we reformulated the density computation to operate primarily on the logarithmic scale.
@@
5.3 Model adequacy and comparison
@@
Table 1: Comparison of LOOIC between AL, GAL, and cGAL models across quantiles. Additionally, the table includes p-values for the scaled residual diagnostics: a Kolmogorov-Smirnov test (uniformity), a dispersion test (under-/overdispersion), and a binomial test for outliers. In each test, higher p-values (e.g., above 0.05) indicate that the data do not deviate substantially from model assumptions, suggesting a better fit.
@@
    $p$-value from the Kolmogorov–Smirnov test (uniformity).
@@ 5.4 Quantile fits
Figure 6 shows posterior estimates and 95% HPD intervals for $\beta_{1}$ to $\beta_{5}$ across five quantiles, while Web LABEL:tab:POST_SUMMARY summarizes parameter estimates and 95% HPD intervals for all models. From Figure 6, we observe the following:

### v2

@@ Abstract
Keywords: Bayesian; Generalized asymmetric Laplace distribution; Mixture distribution; Mixed-effects models; Robust quantile regression; Viral load
Introduction
Mixed-effects quantile regression models have become increasingly important in statistical analyses because they account for the correlation structure inherent in repeated or clustered measurements, thus facilitating more accurate inference of the conditional distribution across various quantiles (Koenker and Hallock, 2001; Geraci and Bottai, 2007). Under a standard mixed-effects framework, this approach is well-suited to heteroscedastic settings, where variability can differ substantially across individuals or groups. In longitudinal clinical trials, for example, certain treatments may have a stronger effect on patients in the lower or upper tails of the response distribution than those near the mean. By capturing these diverse responses, quantile-based methods reveal treatment effects that mean-focused analyses might overlook. The asymmetric Laplace (AL) distribution has long been the favored choice for parametric mixed-effects quantile regression because its fixed skewness parameter directly targets the quantile of interest (Geraci and Bottai, 2007).
@@ Introduction
To overcome the AL distribution’s forced skewness constraint, Yan and Kottas (2017) proposed the generalized AL (GAL) distribution, which gives greater flexibility for asymmetric and heavy-tailed data. Yu and Yu (2023) further extended this approach to mixed-effects models in a quantile regression framework. By adding an estimable shape parameter, the GAL distribution accommodates a wider range of skewness and kurtosis, making it more adaptable than the AL in many settings.
@@
However, while the GAL distribution is more flexible, it can still be vulnerable to extreme outliers, especially in heavy-tailed data. By construction, its location parameter is the $p_{0}$th quantile. When $p_{0}$ is close to 0.5, both tails remain relatively heavy. But as $p_{0}$ moves away from 0.5 (i.e., $p_{0}\to 0^{+}$ or $p_{0}\to 1^{-}$), the skewness parameter “lightens” one tail while weighting the other more heavily, leaving the lighter-tail side more susceptible to outliers. To address this, we propose a more robust approach: the contaminated GAL (cGAL) distribution. Constructed as a mixture of two GAL components, the cGAL inflates the variance of the second component to account for extreme observations, effectively mitigating outlier sensitivity. Crucially, it retains the property that the location parameter aligns with the targeted quantile, preserving interpretability while reducing sensitivity to skewness and outliers.
@@
Other relevant work includes Barata (2021), which examined dynamic quantile regression using the GAL distribution in Bayesian settings, though without extending it to mixed models or multivariate data with cluster-level random effects. Florez et al. (2022) applied quantile regression to multivariate and longitudinal data with the multivariate generalized hyperbolic distribution, which effectively models complex dependence but omits random effects and relies on the Bessel function, complicating computation. Wichitaksorn et al. (2014) explored flexible Bayesian quantile regression via scale mixtures of truncated normals. However, akin to AL, their methods lack asymmetry at middle quantiles and produce lighter tails at extremes, increasing outlier sensitivity. Moreover, these approaches do not extend to random effects. Bernardi et al. (2018) employed the skew exponential power distribution for quantile regression, yielding more flexible tails than the AL but applying it only to fixed-effects models, thus excluding random and mixed model contexts.
@@
In this paper, we adopt a Bayesian approach to estimate parameters of the mixed-effects cGAL model with JAGS (Plummer, 2003). This approach enables full posterior inference and does not rely on asymptotic approximations. Full posterior inference can be especially beneficial for smaller or more complex models. Because JAGS supports flexible, user-defined model specifications, our conceptual approach can be translated directly into code, which conveniently facilitates the implementation of the proposed mixed-effects quantile regression model.
@@
ACTG 315 HIV clinical trial
The ACTG 315 dataset, available in the ushr package in R (Morris et al., 2020), includes longitudinal measurements of HIV viral load (log10 RNA copies/mL) over time. It features data on 46 patients, with the longest measurement recorded on Day 196 after baseline (Day 0). Models of HIV viral load dynamics typically show a biphasic decline: an initial rapid reduction in viral load due to effective antiretroviral therapy, followed by a slower stabilization phase as the virus is further suppressed (Perelson et al., 1996). Traditionally, many analyses (e.g., generalized linear models) have concentrated on the mean response, giving insights into average patient trajectories. Yet such approaches often overlook the considerable variability among patients, as shown in Figure 1. Some patients experience sharper declines, while others have irregular fluctuations or slower viral suppression, indicating outliers or subgroups with distinct response patterns (Nelson and Perelson, 2002). A more flexible framework, such as quantile regression, is therefore needed to capture the full distribution of viral load trajectories and better understand HIV progression across diverse patient subgroups.
@@
Laplace distributions for quantile regression
@@
Asymmetric Laplace distribution
@@
The AL distribution can also be viewed as a normal-exponential mixture (e.g., Kotz et al., 2001):
@@
Generalized asymmetric Laplace distribution
Next, consider the GAL distribution suggested by Yan and Kottas (2017), which generalizes the AL. As with the AL distribution, we fix $\mu$ to be the $p_{0}$th quantile so that
@@
The GAL distribution can also be expressed as a triple mixture of normal, exponential, and truncated normal variables (Yan and Kottas, 2017):
@@
Contaminated generalized asymmetric Laplace distribution
To increase the robustness of a quantile model to outlying values, we introduce the cGAL distribution, formed as a mixture of two GAL distributions. Both components share the same location $\mu$ but differ in scale parameters: $\sigma$ for the main component and $\tau_{0}\sigma$ for the “contaminated” component, where $\tau_{0}>1$. Let $\alpha\in\left(0,1\right)$ be the mixture weight. The cGAL PDF is:
@@
By inflating the scale in the second component, the cGAL distribution can better accommodate outliers. This two-component, scale-inflated mixture approach closely resembles the well-known contaminated normal distribution (Tukey, 1960; Huber, 1992).
@@
Figure 3 presents the sample kurtosis and skewness for the GAL and cGAL ($\tau_{0}=10$, $\alpha=0.5$) distributions as functions of the parameter $\gamma$ to illustrate the impact of contamination in the cGAL distribution on both the tails and asymmetry of the distribution in comparison to the GAL model for two quantiles ($p_{0}\in\left\{0.1,0.5\right\}$). The left and right L-statistic kurtosis (Fiori and Beltrami, 2014) are displayed alongside both distributions’ overall kurtosis and skewness. Details on the calculation of the L-statistic kurtosis can be found in Appendix A. For instance, the GAL distribution exhibits lighter-than-normal left tails at the lower decile ($p_{0}=0.1$). However, with the introduction of contamination in the cGAL distribution, the left tail becomes heavier, making it more robust to outliers, particularly for modeling at the lower decile ($p_{0}=0.1$).
@@
Figure 3: Kurtosis and skewness for GAL and cGAL ($\tau_{0}=10$, $\alpha=0.5$) over $\gamma$. Left and right L-statistic kurtosis (Fiori and Beltrami, 2014) are shown in blue and red, respectively. A dotted black line marks the normal-reference kurtosis of $0.4142$. The overall kurtosis (green) and skewness (black) for GAL ($\tau_{0}=1$, solid) vs. cGAL ($\tau_{0}=10$, dashed) highlight cGAL’s heavier tails and increased robustness. A dotted line at 00 on the skewness plot indicates symmetry.
Bayesian mixed-effects quantile regression model
@@
where $\mu_{ij}\left(p_{0}\right)$ is the conditional $p_{0}^{\text{th}}$ quantile of $y_{ij}$ given the covariates and $\epsilon_{ij}\left(p_{0}\right)$ is the error term such that the $p_{0}^{\text{th}}$ quantile of $\epsilon_{ij}\left(p_{0}\right)$ is zero. Thus, the distribution of $\epsilon_{ij}\left(p_{0}\right)$ is centered at zero in the “$\tau$-quantile” sense.
@@
The mixed-effects quantile model is formulated under the three different distributional frameworks in Section 3: the AL, GAL, and cGAL distributions. In the AL model, $y_{ij}\sim f_{\text{AL}}\left(y_{ij}\left|\mu_{ij}\left(p_{0}\right),\sigma,p_{0}\right.\right)$. The GAL model introduces a skewness parameter, giving $y_{ij}\sim f_{\text{GAL}}\left(y_{ij}\left|\mu_{ij}\left(p_{0}\right),\sigma,\gamma,p_{0}\right.\right)$. Lastly, the cGAL model incorporates an additional contamination parameter to handle outliers, leading to $y_{ij}\sim f_{\text{cGAL}}\left(y_{ij}\left|\mu_{ij}\left(p_{0}\right),\sigma,\gamma,\alpha,p_{0}\right.\right)$.
In the cGAL setting, “outliers” or extreme data points are partially explained by the contaminated component.
For the fixed effects $\bm{\beta}$, we assume a multivariate normal prior with mean $\bm{0}$ and covariance matrix $s_{\beta}^{2}\bm{I}_{p}$:
${\bm{\beta}}\sim\text{MVN}\left({\bm{0}},s^{2}_{\beta}{\bm{I}}_{p}\right),$ (18)
@@
$\sigma\sim t_{+}\left(0,s_{\sigma},\nu_{\sigma}\right),$ (19)
@@
indicating a $t$-distribution with $\nu_{\sigma}$ degrees of freedom, centered at 00, scale $s_{\sigma}$, and truncated to be positive.
For the skewness parameter, we introduce an auxiliary proportion $B\in\left(0,1\right)$ with a beta prior and then rescale it to the admissible interval, $\left(L_{\gamma},U_{\gamma}\right)$:
$B\sim\mathrm{Beta}\left(a_{\gamma},b_{\gamma}\right),\qquad\gamma=L_{\gamma}+\left(U_{\gamma}-L_{\gamma}\right)B,$ (20)
where $E\left[\gamma\right]=a_{\gamma}/\left(a_{\gamma}+b_{\gamma}\right)$ and variance
$\operatorname{Var}\left[\gamma\right]=\dfrac{a_{\gamma}b_{\gamma}}{\left(a_{\gamma}+b_{\gamma}\right)^{2}\left(a_{\gamma}+b_{\gamma}+1\right)}.$ (21)
Setting $a_{\gamma}=b_{\gamma}=1$ reproduces a uniform prior on $\gamma$; alternative choices can up- or down-weight the ends of the interval.
@@
We specify the matrix-generalized half-$t$ (MGH-$t$) prior distribution of Huang and Wand (2013) for the inverse covariance matrix $\bm{\Sigma}^{-1}$ as a hierarchical mixture of a Wishart distribution and gamma distributions for the diagonal elements of a scale matrix, $\bm{\Psi}$. Specifically, the diagonal elements of $\bm{\Psi}$, denoted as $\psi_{ii}$, follow independent gamma distributions with shape and scale parameters set to $0.5$ and $1/A_{\psi}^{2}$, respectively: $\psi_{ii}\sim\text{Gamma}\left(0.5,1/A_{\psi}^{2}\right)$. The off-diagonal elements $\psi_{ij}$ for $i\neq j$ are set to zero. The inverse covariance matrix, $\bm{\Sigma}^{-1}$, follows a Wishart distribution with inverse scale matrix $2\nu_{\Sigma}\bm{\Psi}$ and degrees of freedom equal to $\nu_{\Sigma}+d-1$. We write the MGH-$t$ prior for $\bm{\Sigma}^{-1}$, compactly, as:
$\bm{\Sigma}^{-1}\sim\mathrm{MGH}\text{-}t\left(A_{\psi},\nu_{\Sigma}\right).$ (22)
This hierarchical formulation corresponds to specifying a $t_{+}\left(0,A_{\psi},\nu_{\Sigma}\right)$ prior for the standard deviation terms of $\bm{\Sigma}$.
For the contamination weight, we assign a beta prior
$\alpha\sim\text{Beta}\left(a_{\alpha},b_{\alpha}\right),$ (23)
Choosing $a_{\alpha}\ll b_{\alpha}$ encodes the prior belief that only a small fraction of observations belong to the contaminated (heavier-tailed) component, while still allowing the likelihood to pull $\alpha$ upward when the data warrant it.
Throughout, we treat $\tau_{0}$ as a tuning constant; any sufficiently large value provides heavy-tail protection without altering the model’s quantile-centering property.
The quantities $s_{\beta}^{2}$, $s_{\sigma}$, $\nu_{\sigma}$, $A_{\psi}$, $\nu_{\Sigma}$, $a_{\gamma}$, $b_{\gamma}$, $a_{\alpha}$, $b_{\alpha}$, and the scale-inflation constant $\tau_{0}$ can be chosen to reflect substantive prior beliefs or pragmatic weak-information defaults. Section 5 specifies the values adopted in the HIV analysis and presents sensitivity checks under markedly flatter alternatives.
Application to HIV dataset
Nonlinear mixed-effects model
@@
The NLME model we explore is a specific case of the general mixed-effects quantile regression framework introduced in Section 4, tailored for modeling the HIV viral load in the ACTG 315 dataset. In this case, the model is nonlinear, specified as a biphasic (two-phase) exponential decay process that captures both short-term and long-term dynamics of HIV viral load (Perelson et al., 1996). We model the observed viral load for patient $i$ at time $t_{ij}$ as a combination of two exponential decay processes to explain the biphasic viral decline. The model accounts for both short-term and long-term decay components and includes random effects to capture patient-specific variability.
We model the observed viral load $y_{ij}$ as follows (Wu and Ding, 1999):
@@
The random effects vector $\bm{b}_{i}=\left(b_{i1},b_{i2},b_{i3},b_{i4}\right)^{\top}$ is assumed to follow a multivariate normal distribution as in Equation (17).
@@
The above NLME model, based on the GAL distribution, was also applied to the ACTG 315 dataset by Yu and Yu (2023). However, they implemented a custom sampler instead of JAGS and adopted different prior specifications. For instance, they employed a Wishart prior for $\bm{\Sigma}$, whereas we use the matrix-generalized half-$t$ distribution.
Model implementation, adequacy, and comparison
@@
We adopt the following weakly informative specifications:
$\bm{\beta}$ $\sim\mathrm{MVN}\left(\mathbf{0},1000\mathbf{I}_{p}\right),$ (27)
$\sigma$ $\sim t_{+}\left(0,\sqrt{10},3\right),$
$B$ $\sim\mathrm{Beta}\left(1,1\right),$
$\gamma$ $=L_{\gamma}+(U_{\gamma}-L_{\gamma})B,$
$\bm{\Sigma}^{-1}$ $\sim\mathrm{MGH}\text{-}t\left(A_{\psi}=50,\nu_{\Sigma}=2\right),$
$\alpha$ $\sim\mathrm{Beta}\left(1,9\right),$
$\tau_{0}$ $=10.$
Here $\bm{\beta}=\left(\beta_{1},\dots,\beta_{5}\right)^{\top}$ ($p=5$) and $\bm{\Sigma}$ is a $d\times d$ random-effects covariance matrix with $d=4$; thus the MGH-$t$ construction uses $\nu_{\Sigma}+d-1=5$ degrees of freedom in the implied Wishart step, conditional on the diagonal scale matrix $\bm{\Psi}$.
The prior for $\bm{\beta}$ implies the 95% prior region covers approximately ±$\pm 60$ log10 RNA copies/mL, which exceeds any clinically plausible range. Therefore, the likelihood dominates inference.
The choice of $\alpha\sim\mathrm{Beta}\left(1,9\right)$ implies an expected value of $E\left[\alpha\right]=0.10$ and results in a strongly right-skewed distribution. This prior reflects our expectation that only a small fraction of observations arise from the contaminated component, while still permitting the data to suggest larger values of $\alpha$ if necessary. Similarly, setting $\tau_{0}=10$ increases the scale of the contaminated component tenfold, large enough to capture heavy-tailed outliers without dominating the main GAL component.
All models were implemented and run in JAGS (Plummer, 2003) through the runjags package in R (Denwood, 2016). The Bayesian models were executed using Markov chain Monte Carlo (MCMC) sampling, with multiple chains initiated to ensure a thorough exploration of the posterior distribution. Convergence diagnostics were checked for each model using the potential scale reduction factor ($\hat{R}$), with values below 1.05 indicating acceptable convergence across chains.
For the AL and GAL models, we utilized their mixture representations in Equations (5) and (11) to streamline the MCMC sampling process. In contrast, for the cGAL model, we used the “ones” trick to define its custom likelihood in JAGS. This method uses binary indicator variables that follow a Bernoulli distribution. The probabilities are proportional to the likelihood, scaled by a constant to ensure they remain valid. Compared to the AL and GAL models’ mixture representation, this approach increases computational demands.
@@
The Kullback-Leibler (K-L) divergence is employed to assess the adequacy of the three models to identify influential observations and outliers (Wang and Luo, 2016; Tomazella et al., 2021), as outlined in Appendix B.
We conducted a simulation-based residual analysis using the R package DHARMa (Hartig, 2021a, b) to evaluate the adequacy of the three models. This method compares the observed data to the posterior predictive distribution, conditioned on the random effects for each observation, thereby generating scaled residuals between 0 and 1. We used several tests provided by DHARMa to assess these scaled residuals statistically. The Kolmogorov-Smirnov test evaluated the uniformity of the residual distribution; a parametric dispersion test checked for overdispersion or underdispersion, and a binomial test flagged observations significantly deviating from model expectations. We report p-values from these tests to determine each model’s adequacy of fit.
The loo package (Vehtari et al., 2024) in R was used to perform leave-one-out (LOO) cross-validation to compare the models’ predictive performance. To ensure numerical stability in the computation of the GAL and cGAL density functions, especially when dealing with exponentials of large magnitudes, we adopted a numerically stable implementation of the log-density function. The original GAL density function involves exponential and CDF terms that can cause overflow or underflow issues. To mitigate this, we reformulated the density computation to operate primarily on the logarithmic scale.
@@
Model adequacy and comparison
@@
Table 1: Comparison of LOOIC between AL, GAL, and cGAL models across quantiles. LOOIC is derived from full LOO cross-validation; it equals 2×-2$\times$- 2 × the sum of the log-predictive densities of each observation given a model refit to the remaining data. Additionally, the table includes p-values for the scaled residual diagnostics: a Kolmogorov-Smirnov test (uniformity), a dispersion test (under-/overdispersion), and a binomial test for outliers. In each test, higher p-values (e.g., above 0.05) indicate that the data do not deviate substantially from model assumptions, suggesting a better fit.
@@
    $p$-value from the Kolmogorov-Smirnov test (uniformity).
@@
Figure 6 shows posterior estimates and 95% HPD intervals for $\beta_{1}$ to $\beta_{5}$ across five quantiles, while Web LABEL:tab:POST_SUMMARY summarizes parameter estimates and 95% HPD intervals for all models. Web LABEL:tab:POST_SUMMARY also reports results under markedly flatter priors; these are discussed in Section 5.5. From Figure 6, we observe the following: